A drunkard is walking along a straight road. He takes 5 steps forward and 3 steps backward, followed by 5 steps forward and 3 steps backward and so on. Each step is one meter long and takes one second. There is a pit on the road 11 meter away from the starting point then after how much time drunkard will fall into the pit?

Grade:12th pass

5 Answers

Vikas TU
7 years ago
For  8 total steps it takes 8 seconds,
and moves in total 2 steps forward (2 m) 
Therfore,
3*2 m = 6 m 
At 3rd cycle Time Elapsed  24 s 
After walking 5 m forward, he takes 5 s more.
It becomes 11 m. 
Time = 24 + 5 = 29m
Saurav
6 years ago
2m-8sec4m-16secAnd so on,,, 4*6=24 And it will need 5sec more to fell down in yard so 24+5=29sec ,,,thanks. ,,,,,,
Krrish
6 years ago
Distance covered with 1 step = 1 m
Time taken = 1 s
Time taken to move first 5 m forward = 5 s
Time taken to move 3 m backward = 3 s
Net distance covered = 5 – 3 = 2 m
Net time taken to cover 2 m = 8 s
Drunkard covers 2 m in 8 s.
Drunkard covered 4 m in 16 s.
Drunkard covered 6 m in 24 s.
Drunkard covered 8 m in 32 s.
In the next 5 s, the drunkard will cover a distance of 5 m and a total distance of 13 m and falls into the pit.
Net time taken by the drunkard to cover 13 m = 32 + 5 = 37 s
Ajeet Tiwari
3 years ago
hello students

the person moves 5 steps forward and 3 steps backward in 8sec after that he is 2m away from an initial position
x= 5-3=2m in 8 sec
after 24sec he will 6m away from an initial position
further he moves 5m forward then he falls in a pit which is 11m away from the initial position.
total time T=24+5=29sec.

hope it helps
thankyou
Yash Chourasiya
3 years ago
Dear Student

The person moves 5 steps forward and 3 steps backward in 8sec after that he is 2m away from an initial position
x = 5-3 = 2m in 8 sec
after 24sec he will 6m away from an initial position
further he moves 5m forward then he falls in a pit which is 11m away from the initial position.
total time T = 24 + 5 = 29sec.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya